OpenStudy (samigupta8):

Two particles of mass m and 2m are fixed to a thin uniform circular hoop of mass m ,radius 30/22 cm at its diametrically opposite ends(topmost and bottommost points) .The velocity of centre of hoop is Vo when it is in pure rolling. the maximum value of Vo in m/s so that hoop does not lose its contact with ground is

OpenStudy (samigupta8):

@michele_laino

OpenStudy (irishboy123):

contact = 0 at: \(\dfrac{2m v_0^2}{r} = mg + \dfrac{m v_0^2}{r}\)

OpenStudy (irishboy123):